# Squeezing information from radio surveys to probe the primordial Universe

### v1

@@ IV Results
Here we only consider a squeezed subset of these configurations, i.e. those that satisfy $k_{1}\approx k_{2}\geq 3\,k_{\rm 3}$, since they contain most of the signal for the local PNG. Note, that the bimodal bispectrum can benefit all triangle shapes (see the Appendix). The minimum value is $k_{\rm min}=k_{\rm f}$, which is the largest scale probed in the redshift bin by the survey, and the maximum value $k_{\rm max}$ corresponds to the smallest scale where the theoretical model is reliable. We use a conservative choice $k_{\rm max}(z)=0.75\,k_{\rm NL}(z)$ and an optimistic one $k_{\rm max}(z)=k_{\rm NL}(z)$, where $k_{\text{NL}}$ is given by the inverse square root of the one-dimensional velocity dispersion [61]. The analysis is thus confined within the perturbative regime, where the tree-level model offers good agreement with the numerical results [72, 73, 74, 75, 76]. The tree-level description is used to model the expected value of the bimodal estimator [Section III], including the observational windows for the two operating modes of the instrument. Any $\bm{k}$-mode outside of the window will result in the exclusion of the triangle from the Fisher matrix calculation.
Figure 2: Marginalized forecasts of $\sigma(f_{\mathrm{NL}}^{\mathrm{loc}})$ from the bimodal bispectrum (SKA $\times$ HIRAX) with and without foreground cut, for pessimistic and optimistic $k_{\rm max}$. Forecasts from each survey and from the summed signal without cross correlation (SKA + HIRAX) are also shown. Purple line is the Planck 2018 constraint [77].
The forecast error on $\theta_{\alpha}$ is then given by $\sigma(\theta_{\alpha})=(\tilde{F}_{\alpha\alpha}^{-1})^{1/2}$, where $\tilde{F}_{\alpha\beta}=\Sigma_{i}F_{\alpha\beta}(z_{i})$ encompasses the total information from all redshift bins considered. All the parameters that control the model are considered free and are entries of the Fisher matrix. These include the $\Lambda$CDM cosmology parameters and the HI IM bias parameters [9, 61, 58, 78] (see the Appendix, and [79] for further details on the Fisher formalism).
All work on constraining $f_{\mathrm{NL}}^{\mathrm{loc}}$ is affected by a degeneracy between $f_{\mathrm{NL}}^{\mathrm{loc}}$ and $b_{\phi}$ (see e.g. [80, 51]), where $b_{\phi}$ is the local PNG bias parameter [48, 81]. This compromises the ability of the LSS power spectrum and bispectrum to tightly constrain $f_{\mathrm{NL}}^{\mathrm{loc}}$. PNG bias depends on the halo formation history [82], i.e., on properties beyond total mass. If it is accurately modeled and constrained, e.g. with hydrodynamical simulations, this degeneracy is broken and the power of the LSS bispectrum to constrain $f_{\mathrm{NL}}^{\mathrm{loc}}$ is completely restored [52]. Our focus here is on the relative improvement from a new bimodal synergy, compared to the sum of the standard bispectrum information from 2 surveys – therefore, our results are not affected by the model for $b_{\phi}$, and we use the simplest model that breaks the degeneracy [44, 46, 45].
After marginalization over the free parameters of the model [58, 78] (see also the Appendix), we retrieve the final forecasts on $f_{\mathrm{NL}}^{\mathrm{loc}}$ from the bimodal bispectrum. The two surveys that we consider are Band 1 of SKA-MID [29], which is a radio telescope in SD mode and has a redshift range of $z=0.35-3.05$, and the final version of HIRAX [30] in IF mode with $z=0.75-2.55$. In the analysis, we consider only the signal from the overlapping range ($z=0.75-2.55$) and sky coverage ($15,000$ $\rm deg^{2}$). For details of the surveys considered and the specifics of the instrumental noise, see [78]. Note that the same survey can be used, operating in SD and IF mode, to produce the bimodal synergy. However, telescope design is usually optimized for one of the modes and is not effective in the other.
The forecasts on $f_{\mathrm{NL}}^{\mathrm{loc}}$ from the synergy between SKA and HIRAX are presented in Fig. 2, for the pessimistic and optimistic $k_{\rm max}$ cases. The results when removing the foreground cut, i.e. when assuming perfect foreground cleaning and signal reconstruction, are shown in order to illustrate the limiting capabilities of the surveys. We also present the forecasts from each individual survey, as well as their summed signal (SKA+HIRAX) without cross-correlation, for the same redshift range and squeezed configurations as in the bimodal case. This allows for a fair comparison and an assessment of the gain from the proposed method. The results indicate that the bimodal bispectrum provides a significant improvement on $f_{\mathrm{NL}}^{\mathrm{loc}}$ constraints, more than a factor of $3$, compared to those from each individual survey, as well as their summed signal (a factor of $\sim 3$).
The constraints are competitive with current CMB bounds, exhibiting a factor of $\sim 2$ improvement. The bimodal bispectrum achieves this by only using a small subset, $\mathcal{O}(10^{2}-10^{3})$ triangles, of the total number of formed triangles. This means that the forecasts could improve further once all the triangles are considered, even achieving $\sigma(f_{\mathrm{NL}}^{\mathrm{loc}})\lesssim 1$, as in the case of dedicated surveys (e.g. SPHEREx [83] or PUMA [84]). In addition, by including the single-survey information from the non-overlapping redshifts and sky area, the constraints would improve even further and utilise to the fullest the available signal in the SKA-HIRAX synergy.
V Conclusions
The MeerKLASS pilot survey [85] on the MeerKAT precursor telescope of SKA-MID, has provided HI IM maps both in SD and IF mode. This has allowed the detection of the SD power spectrum in cross correlation with galaxies [86] and the auto-power in IF mode [87]. The measurements in SD mode will see a major advance when the SKA is operating, while the early IF mode measurements will be vastly improved when HIRAX takes data.
We have shown how to combine these observations in a novel way to give significantly stronger constraints on $f_{\mathrm{NL}}^{\mathrm{loc}}$ than was previously thought possible. In fact, this method will improve on the Planck results by around a factor of 2, using only a few thousand triangles per redshift bin. The bimodal bispectrum improves constraints relative to the sum of the individual surveys by a factor of $\sim 4$, which is equivalent to a much larger SD or IF array. It is in fact the only way to get any strong constraints on local PNG from the HI intensity bispectrum.
We highlight the point that a substantial part of the constraining power comes from the IF mode signal. Therefore, the results will improve even further if the IF instrumental noise can be reduced. By contrast, the method is not so sensitive to the SD mode noise (see Fig. 4). Instead, the SD signal is sensitive to the loss of very large scales $k_{\parallel}<k_{\parallel{\rm min}}$ (see Figs. 3 and 5), so that further improvement requires advances in foreground cleaning and long-wavelength $k_{\parallel}$-mode reconstruction.
Long-wavelength modes provide a set of very squeezed triangles that carry a significant amount of signal on local PNG. On the other hand, increasing the number of squeezed triangles by changing $k_{\rm max}$ (i.e. the pessimistic and optimistic cases) provides a much less pronounced gain in the $f_{\mathrm{NL}}^{\mathrm{loc}}$ constraints ($\sim 20\%$). Notably, the constraints with an optimistic $k_{\rm max}$ are very close ($\sim 15\%$ difference) to those provided by the pessimistic case, in which no radial foregrounds are considered (see Fig. 2). This indicates that the local PNG signal lost to foregrounds can be partially recovered, by venturing further into the intermediate regime. Smaller scales have a significant amount of information on PNG, which requires alternative methods to the analytical modelling, e.g. simulated-based inference [17, 18, 19, 20]. We leave this for future work.
The bimodal bispectrum cross-correlates SD large scales with smaller-scale IF modes. We do not expect foregrounds or instrumental systematics to show correlations between long- and short-wavelength $k$-modes. This is another advantage of the bimodal synergy. The advantages of the bimodal bispectrum are not limited to local PNG studies. Any observable that has signal peaking on the squeezed triangles can benefit from the proposed synergy. Moreover, including all shapes of triangles into the bimodal analysis (see Fig. 6), still maintains a significant gain in the constraining power.
Acknowledgements DK, RM and SS acknowledge support for this work from the University of Missouri South African Education Program. RM is supported by the South African Radio Observatory and National Research Foundation (grant no. 75415). SS acknowledges support for this work from NSF-2219212. SS is supported in part by World Premier International Research Center Initiative, MEXT, Japan. SC acknowledges support from the Italian Ministry of University and Research, PRIN 2022 ‘EXSKALIBUR – Euclid-Cross-SKA: Likelihood Inference Building for Universe Research’, from the Italian Ministry of Foreign Affairs and International Cooperation (grant no. ZA23GR03), and from the European Union – Next Generation EU. JF acknowledges support of Fundação para a Ciência e a Tecnologia through the Investigador FCT Contract No. 2020.02633.CEECIND/CP1631/CT0002, the FCT project PTDC/FIS-AST/0054/2021, and the research grants UIDB/04434/2020 and UIDP/04434/2020.

### v2

@@ Appendix A HI IM model and forecast details
The full set of parameters consists of 5 cosmological parameters, the primordial PNG amplitude and 11 further parameters in each redshift bin $i$:
$\bm{\theta}(z_{i})=$ $\Big{\{}\Omega_{\rm b},\Omega_{\rm c},h,A_{\rm s},n_{\rm s},f_{\mathrm{NL}}^{\mathrm{loc}};$ (A.5)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ D_{A}(z_{i}),H(z_{i}),f(z_{i}),b_{1}(z_{i}),b_{2}(z_{i}),b_{s^{2}}(z_{i}),\sigma_{P}(z_{i}),\sigma_{B}(z_{i}),P_{\varepsilon}(z_{i}),P_{\varepsilon\varepsilon_{\delta}}(z_{i}),B_{\varepsilon}(z_{i})\Big{\}}.$
For the 5 cosmological parameters we add prior information from the observation of CMB performed by Planck [97]. We assume redshift bins are independent, so that e.g. $\partial P(\bm{k},z_{i})/\partial\theta_{\alpha}(z_{j})=0$ for $i\neq j$. All the above parameters are considered unknown and are marginalized over in the Fisher matrix formalism. Note that we also consider the anisotropies induced in the observed galaxy clustering by the Alcock-Paczynski effect [98], which occurs when the fiducial cosmology, used to convert the observed angular coordinates and redshifts to physical distances, differs from the true one [99, 100]. The fiducial values used for the cosmological parameters are those measured by Planck [97] and are presented for convenience in Table˜2 . The redshift-dependent expressions for the HI bias parameters are provided in [75]. The PNG bias parameters, $b_{\phi}(b_{1})$ and $b_{\phi\delta}(b_{1})$, are considered a function of the linear bias, i.e. $b_{\phi}(b_{1})$ and $b_{\phi\delta}(b_{1})$, using the ‘universality relation’ [33], where the expressions are presented in e.g. [57]. These PNG bias parameters are considered fixed throughout the analysis.
[…]
As emphasized in the main text, accessing sufficiently long-wavelength $k$-modes with the SD survey is important for the proposed method to deliver competitive $f_{\mathrm{NL}}^{\mathrm{loc}}$ constraints, irrespective of whether radial foregrounds are mitigated. This depends on how effectively the SD survey can probe large scales. To investigate this further, we generate $f_{\mathrm{NL}}^{\mathrm{loc}}$ forecasts using the methodology described in the main text, varying key instrumental parameters that influence the ability of the SD survey (i.e., SKA) to probe large scales. Specifically, we consider the total integration time ($t_{\rm tot}$) and the number of dishes ($N_{\rm dish}$), both of which impact the instrumental noise. Increasing either parameter leads to reduced instrumental noise levels and, consequently, improved precision in $f_{\mathrm{NL}}^{\mathrm{loc}}$ measurements.
[…]
The forecast $\sigma(f_{\mathrm{NL}}^{\mathrm{loc}})$, normalized to the fiducial SD survey specifications used to derive the main results (Fig.˜2), is shown in Fig.˜4 as a function of $t_{\rm tot}$ (left panel) and $N_{\rm dish}$ (right panel). Decreasing by an order of magnitude the capabilities of the SD mode survey (e.g. considering only 20 dishes out of 200 or 1000 hrs instead of 10000 hrs), degrades the $f_{\mathrm{NL}}^{\mathrm{loc}}$ constraint by only $\sim 10\%$. This indicates that the effect of the instrumental noise of the SD mode survey is small, since it is still able to efficiently probe the large scales.
We use as an example a variation of the SD survey, ‘SKA/2’, where the integrated time $t_{\rm tot}$ and the number of dishes $N_{\rm dish}$ are reduced by half. The bimodal bispectrum $f_{\mathrm{NL}}^{\mathrm{loc}}$ forecasts are presented in Fig.˜5 (SKA/2 $\times$ HIRAX) by following the process described in the main text. Moreover, the idealized case where the radial $k$-modes lost to foregrounds are reconstructed (i.e. no $k_{\parallel{\rm min}}$ cut), is also shown in dashed lines. The results show that reducing the capabilities of the SD survey by half, has minimal effect on the bimodal constraints on $f_{\mathrm{NL}}^{\mathrm{loc}}$. This is true for both values of $k_{\rm max}$ (‘conservative’ and ‘Optimistic’) and for the cases where the modes $k_{\parallel}<k_{\parallel{\rm min}}$ are included (dashed lines) and excluded (solid lines) from the analysis. The tightest constraint, $\sigma(f_{\mathrm{NL}}^{\mathrm{loc}})\sim 2$, corresponds to the optimistic $k_{\rm max}$ and the case with no $k_{\parallel{\rm min}}$ cut.
The advantages of the bimodal bispectrum are not limited to squeezed configurations and other shapes can be utilised. In Fig.˜6 we show forecasts on $f_{\mathrm{NL}}^{\mathrm{loc}}$ when all triangle configurations are used and a redshift independent $k_{\rm max}=0.2\;\text{$h$/Mpc}$ is considered. Note that the complete bimodal bispectrum covariance is used in these forecasts, since the approximate expression for the non-Gaussian part [Section˜3] is accurate, for the scales considered, in the case of all shapes of triangles [82]. The bimodal bispectrum still provides a significant improvement over the constraints from each individual survey (a factor of $\sim 2$), while it exhibits a $\sim 30\%$ improvement over their summed signal. Pushing $k_{\rm max}$ to smaller values would enhance these differences even further, as shown in the main text.
Figure 6: Same as Fig.˜2, but for all triangle shapes, not only squeezed, and with a fixed $k_{\rm max}(z)=0.2\;\text{$h$/Mpc}$.